find 2 numbers with the mean of 14 and a difference of 4
Answer by stanbon(75887)

find 2 numbers with the mean of 14 and a difference of 4
-------------
Mean: (x+y)/2 = 14
Diff: x-y = 4
--------------------
y = x-4
Substitute to get:
(x+x-4) = 28
2x = 32
x = 16
--------
Solve for "y":
x-y = 4
16-y = 4
y = 12
===================
Cheers,
Stan H.